The Family Planning Association in Oxford, England have followed 17,000 women who enrolled in a cohort study between 1988 and 1994. A recent publication compared women currently using oral contraceptives (OCP) with women who were never used or were past users of OCP and reported venous thromboembolism (DVT).

For current users of the OCP, person-years of observation were counted from the time a woman began using the OCP. For never or past users, person-years of observation were counted from the time a woman enrolled in the study.

Woman-years were counted until (1) venous thromboembolism occurred, (2) the woman was lost-to-follow-up, or (3) the time the data were analysed; reported results:

  • Current OCP users IR DVT = 68 per 100,000 person-years
  • Never or past OCP users IR DVT = 8 per 100,000 person-years
  • OCPs with higher dosages of oestrogen IR DVT = 58 per 100,000 person-years
  • OCPs with low dosages of oestrogen IR DVT = 31 per 100,000 person-years

  1. Calculate the RR DVT for current users compared to never or past users

  1. Calculate the RR DVT for high dose users compared to never or past users

  1. Calculate the RR DVT for low dose users compared to never or past user

  1. Calculate the RR DVT for high dose users compared to low dose users
Answer #1

Relative Risk formula

/ ( AA + B) C/CD) RR =

a. The relative risk DVT for current users compared to never or past users

A = the number of women who used oral contraceptives and developed DVT = 68

C = The number of women who never used OCPs but developed DVT = 8

A + B = Total number of persons using OCPS may or may not develop DVT = 100,000

C +D = Total number of women never used may or may not develop DVT = 100,000

RR = (68 / 100,000) / (8/100,000)

RR = 0.00068 / 0.00008

= 8.5

The relative risk for DVT for current users compared to never or past users is 8.5. This shows there is a high risk for OCPS users compared to non-users.

b. Calculating the relative risk DVT for high dose users compared to never or past users.

A = the number of high dose users of estrogen developed DVT= 58

C = The number of women who never used OCPs but developed DVT= 8

A + B = Total number of women Who used higher doses of estrogen may or may not develop DVT = 100,000

C +D = Total number of women never used estrogen May or may not develop DVT = 100,000

RR = (58 /100,000) / (8/100,000)

RR = 0.00058 / 0.00008

= 7.25

The relative risk DVT for higher doses estrogen compared to never or past users of OCPS is 7.25. This shows the use of a higher dose of estrogen is at more risk for DVT.

C. Calculating the relative risk DVT for low dose users of estrogen compared to never or past users.

A = the number of women who use low doses of estrogen developed DVT = 31

C = the number of women who never used OCPs or in the past but developed DVT = 8

A + B = Total number of women who used lower doses of estrogen may or may not develop DVT = 100,000

C + D = Total number of women who never used or used in the past may or may not develop DVT = 100,000

RR = (31 / 100,000) / (8 / 100,000)

RR = 0.00031 / 0.00008

= 3.874

The RR DVT for lower doses of estrogen compared to never or past users of OCPS is 3.874. This shows the lower dose estrogen users are relatively high risk for developing DVT.

d. Calculating the RR DVT of high dose users compared to low dose users

A = The total number of women who used a higher dose of estrogen developed DVT = 58

C = The total number of women who used a lower dose of estrogen developed DVT = 31

A + B = The total number of women who used a higher dose of estrogen may or may not develop DVT = 100,000

C +D = The total number of women who used a lower dose of estrogen may or may not develop DVT = 100,000

RR = (58 /100,000) / (31 / 100,000)

RR = 0.00058 / 0.00031

= 1.87

The RR DVT for a higher dose of estrogen compared to the lower dose of estrogen is 1.87. This shows the higher dose of estrogen users have a greater risk of developing DVT.